OpenStudy (anonymous):

2: P(3 and even) is even 2?

OpenStudy (anonymous):

2, 4, 6 are even. Since there are 3 evens and 6 total P(even) = 3/6 = 1/2 and P(3) = 1/6

OpenStudy (anonymous):

now what do you thing we should do?

OpenStudy (anonymous):

so wait the even = 1/2

OpenStudy (anonymous):

yes so to get the final answer, we multiply p(even) and p(3) 1/2 * 1/6 = 1/12

OpenStudy (anonymous):

agina?

OpenStudy (anonymous):

yep, same answer.

OpenStudy (anonymous):

So if it was P (odd and even) it would be P (odd)=3/6=1/2 P(even)=2/6=1/2 P(odd)*(even)=1/4

OpenStudy (anonymous):

yes :) You got it!

OpenStudy (anonymous):

:D thanks :D